OpenStudy (anonymous):

I have a question #14

OpenStudy (anonymous):

\[f(x)=x^4+2x^3-x^2+ax+b\] where \[f(x)=[Q(x)]^2\] find \[\int\limits\limits_{0}^{1}f(x)dx=?\]

OpenStudy (anonymous):

in fraction form

OpenStudy (anonymous):

btw what is Q(x) ?

OpenStudy (anonymous):

both f(x) and Q(x) are polynomial functions.- this is what the problem says

OpenStudy (anonymous):

ok well i'll translate everything they say in the problem

OpenStudy (anonymous):

define: a and b are real numbers f(x) is a polynomial function where \[f(x)=x^4+2x^3-x^2+ax+b\] if there exist a polynomial function Q(x) where \[f(x)=(Q(x))^2\] find \[\int\limits_{0}^{1} f(x) dx =?\] a. \[\frac{71}{30}\] b. \[\frac{31}{30}\] c. \[\frac{11}{30}\] d. \[\frac{1}{30}\] Hope that there's no mistake in my translation

OpenStudy (anonymous):

this is all we get from the problem

OpenStudy (anonymous):

11/30

OpenStudy (anonymous):

correct ffm plz explain

OpenStudy (anonymous):

You don't know how to solve this one ?

OpenStudy (anonymous):

or testing me ? :P

OpenStudy (anonymous):

both ;P

OpenStudy (anonymous):

Okay Let, \( Q(x)= x^2 + P x + Q\) then \( f(x) = Q^2+2 P Q x+P^2 x^2+2 Q x^2+2 P x^3+x^4 \) after this you just need to compare the coefficients and find \( a \) and \( b \) then ...